Thermal Conductivity Ranking Task

Six objects are placed in a 500^ circ rm F (260^ circ rm C) oven and allowed to reach thermal equilibrium. Eachobject has a mass of 1.0 {rm kg}. The specific heat and thermalconductivity of each substance are denoted by c and k.Rank these objects on the basis of their temperatures when removed from the oven.Rank from largest to smallest. To rank items as equivalent, overlap them.

Wooden cutting board

c=1700J/Kg*Ck=0.1 J/(s*m*C)

Silveringotc=235 J/Kg*C
k=420 J/(s*m*C)

Aluminumpotc=900J/Kg*Ck=220J /(s*m*C)

Ironskilletc=448J/Kg*Ck=80 J/(s*m*C)

Glass casseroledishc=837J/Kg*Ck=0.8 J(s*m*C)

Well donesteakc=3500J/Kg*Ck=0.2 J(s*m*C)
Answer #1

General guidance

Concepts and reason

The main concepts used to solve this problem are specific heat and Thermal conductivity.

Initially, Use the definition of the specific heat. Later, use the expression of the thermal conductivity to identity the rankings. Finally, Use the fact that at equilibrium and object and its surroundings are at the same temperature.

Fundamentals

Thermal conductivity is the property of a material to conduct heat. It is defined as the heat energy transferred per unit time and per unit surface area. The objects with low thermal conductivity has low rate of heat transfer across materials than across materials with high thermal conductivity.

Specific heat is defined as the amount of heat per unit mass required to raise the temperature of a substance by one degree Celsius. A high specific heat of a substance implies that it takes a lot of energy to raise its temperature. The water has high specific heat.

Step-by-step

Step 1 of 2

Rank these objects on the basis of their temperatures when removed from the oven.

The specific heat of a substance is the amount of heat energy needed to raise the temperature of 1.0 kg of the substance by 1\u2218C1^\\circ {\\rm{C}}.

Explanation | Hint for next step

The specific heat of the substance depends on the material of the substance. It\u2019s value of different for different substances.

Step 2 of 2

Thermal equilibrium is a state in which there is no heat flows between two physical systems.

At thermal equilibrium, an object and its surroundings are at the same temperature

So, all the substances in an oven maintain same temperature. Because, at thermal equilibrium the substances don\u2019t depends on its specific heat and thermal conductivity.

Ranking from largest to smallest is as follows:

Ironskillet=welldonesteak=glasscasseroledish=aluminumpot=woodencuttingboard=silveringot\\begin{array}{c}\\\\{\\rm{Iron skillet}} = {\\rm{well done steak}} = {\\rm{glass casserole dish}} = {\\rm{aluminum pot}}\\\\\\\\ = {\\rm{wooden cutting board}} = {\\rm{ silver ingot}}\\\\\\end{array}


Explanation

The rankings are found using the fact that at equilibrium the temperature of the all substances are same.

Answer

Ranking from largest to smallest is as follows:

Ironskillet=welldonesteak=glasscasseroledish=aluminumpot=woodencuttingboard=silveringot\\begin{array}{c}\\\\{\\rm{Iron skillet}} = {\\rm{well done steak}} = {\\rm{glass casserole dish}} = {\\rm{aluminum pot}}\\\\\\\\ = {\\rm{wooden cutting board}} = {\\rm{ silver ingot}}\\\\\\end{array}
Answer #2

\\begin{array}{c}\\\\{\\rm{Iron skillet}} = {\\rm{well done steak}} = {\\rm{glass casserole dish}} = {\\rm{aluminum pot}}\\\\\\\\ = {\\rm{wooden cutting board}} = {\\rm{ silver ingot}}\\\\\\end{array}